OpenStudy (anonymous):

Sally was asked to find two integers such that one is three less than 4 times the other and the sum of the two integers is 42. The integers must be _____. A.14, 28 B.15, 27 C.9, 33 D.-13, 55

OpenStudy (anonymous):

it's C. 9*4= 36 36-3= 33 9+33=42

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Np :)